OpenStudy (anonymous):

divide using synthetic division (5x^3 - 11x^2 -17x +15) / (x-3)

OpenStudy (anonymous):

write out the coefficients 5 -11 -17 15 3 15 12 -15 ________________ 5 4 -5 0

OpenStudy (anonymous):

now a word of explanation. you are dividing by x - 3. list the coefficients and put a 3 down because you are dividing by x - 3. drop the 5 3*5=15 11+15=4 3*4=12 -17+12=-5 3*-5=-15 -15+15=0 you have divided a polynomial of degree 3 by a polynomial of degree 1 which gives a polynomial of degree 2. the coefficients are 5, 4 and -5 and there is no remainder so it divides evenly. that means \[\frac{5x^3-11x^2-17x+15}{x-3}=5x^2+4x-5\] or \[5x^3-11x^2-17x+15=(x-3)(5x^2+4x-5)\]

OpenStudy (anonymous):

that also means 3 is a zero of the original polynomial that is, if \[f(x)=5x^3-11x^2-17x+15\] then \[f(3)=0\] because \[f(x)=(3-3)(5\times 3^2+4\times 3-5)=0\]

OpenStudy (anonymous):

oops typo. i meant f(3)=0, not f(x) = 0. sorry.

OpenStudy (anonymous):

THANK TOY SO MUCH

OpenStudy (anonymous):

YOU* ..LOL

OpenStudy (anonymous):

welcome!